OpenStudy (anonymous):

Choose all of the following equations that have a graph in which the parabola opens downward. More than one answer may be chosen. y = 2x^2 + 2x + 8 y = −4x^2 + 8x − 19 y = 16x^2 + 2x − 4 y = −x^2 − 6x + 10

OpenStudy (lgbasallote):

parabola is a quadratic function...that means the equations with square downward means negative coefficient of x^2

OpenStudy (anonymous):

Oh man.. that word.

OpenStudy (anonymous):

Coefficients.

OpenStudy (anonymous):

throws me off everytime.

Parth (parthkohli):

Coefficient of any variable in a term is the number you are multiplying the variable with :)

OpenStudy (lgbasallote):

coefficient is the number beside x lol

Parth (parthkohli):

For example, the coefficient of x in 4x is 4, and the coefficient of 4 in 4x is x.

OpenStudy (anonymous):

left side or right side?

OpenStudy (lgbasallote):

gtg late for class...you get it though..right @careless850 ??

OpenStudy (anonymous):

I'll never get it. :/

OpenStudy (anonymous):

Thanks though.

OpenStudy (lgbasallote):

the number on the right is the exponent..the left is the coefficient

OpenStudy (anonymous):

Okay so what does coefficients have to do with this again?

OpenStudy (anonymous):

So the answers are B and D?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ahh...

OpenStudy (anonymous):

So if these are - squared then it opens downward. If they are + it's a U.

OpenStudy (anonymous):

Yes. Do you know about maxima and minima?

OpenStudy (anonymous):

Choose all of the following equations that have a vertex that is the minimum point of the parabola. More than one answer may be chosen. y = 2x2 + 2x + 8 y = −16x2 + 2x − 4 y = −x2 − 4x − 3 y = x2 + 2x + 4

OpenStudy (anonymous):

Just got to those. Lol.

OpenStudy (anonymous):

It is directly related. If a quadratic, N, opens down, it has a maxima. If a quadratic, P, opens up, it has a minima. http://www3.wolframalpha.com/Calculate/MSP/MSP30601a1hbiag5d670d4100001hhh30h2fgf00h41?MSPStoreType=image/gif&s=47&w=299&h=130&cdf=RangeControl This is a quadratic that opens downward. Do you see how it has a maxima? In other words, how it has a maximum value? http://www3.wolframalpha.com/Calculate/MSP/MSP22531a1hc11hcci924d300000g2h9ag0b56c518b?MSPStoreType=image/gif&s=45&w=299&h=139&cdf=RangeControl This is a quadratic that opens upward. Do you see how it has a minima? In other words, how it has a minimum value?

OpenStudy (anonymous):

Minimas are wider U ... and maxima are wider upside down u.

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

they go x from -1 to 1.

OpenStudy (anonymous):

Not quite. A quadratic *has* a minima. But it only has a minima if it opens upward. Likewise, a quadratic *has* a maxima. But it only has a maxima if it opens downward. (The pictures are just the typical example.) Can you answer your question now? :)

OpenStudy (anonymous):

A & D.

OpenStudy (anonymous):

Good job! :D